OpenStudy (anonymous):

The Ksp for magnesium hydroxide (Mg(OH)2) is 1.1 x 10-11. What is the molar solubility of the magnesium ion if the pH is 10.0?

OpenStudy (rogue):

pH + pOH = 14 pH = 10 ---> pOH = 4 [OH-] = 10^(-pOH) [OH-] = 10^-4 M \[Mg(OH)_2 (s) \rightarrow Mg^{2+} (aq) + 2OH^- (aq)\]\[K_{sp} = [Mg^{2+}][OH^-]^2\]\[1.1 \times 10^{-11} = [Mg^{2+}] \times (10^{-4})^2\]\[[Mg^{2+}] = \frac {1.1 \times 10^{-11}}{10^{-8}} = 1.1 \times 10^{-3} M\]